### v1

# Learning to generate feasible graphs using graph grammars

Abstract
Generative methods for graphs need to be sufficiently flexible to model complex dependencies between sets of nodes. At the same time, the generated graphs need to satisfy domain-dependent feasibility conditions, that is, they should not violate certain constraints that would make their interpretation impossible within the given application domain (e.g. a molecular graph where an atom has a very large number of chemical bounds). Crucially, constraints can involve not only local but also long-range dependencies: for example, the maximal length of a cycle can be bounded.
Currently, a large class of generative approaches for graphs, such as methods based on artificial neural networks, is based on message passing schemes. These approaches suffer from information ’dilution’ issues that severely limit the maximal range of the dependencies that can be modeled. To address this problem, we propose a generative approach based on the notion of graph grammars. The key novel idea is to introduce a domain-dependent coarsening procedure to provide short-cuts for long-range dependencies.
We show the effectiveness of our proposal in two domains: 1) small drugs and 2) RNA secondary structures. In the first case, we compare the quality of the generated molecular graphs via the Molecular Sets (MOSES) benchmark suite, which evaluates the distance between generated and real molecules, their lipophilicity, synthesizability, and drug-likeness. In the second case, we show that the approach can generate very large graphs (with hundreds of nodes) that are accepted as valid examples for a desired RNA family by the “Infernal” covariance model, a state-of-the-art RNA classifier.
Our implementation is available on github: github.com/fabriziocosta/GraphLearn
1 Introduction
Generative methods for graphs are receiving increased attention due to the variety and importance of application domains that can be naturally modeled using this formalism. Operating on discrete structures however often incurs computational problems due to the combinatorial complexity explosion. One popular domain where these problems are manageable is organic chemistry, where data is plentiful and the size of the discrete label’s alphabet and of the overall molecular graph are small (Polykovskiy et al., 2020). However, larger molecules, such as RNAs or proteins, exhibit an additional source of difficulties, i.e. the presence of complex dependencies between their different constituent parts.
In Costa (2017) the problem of generating elements of a structured domain was framed as the equivalent problem of sampling from a corresponding underlying probability distribution defined over a (learned) class of structures. Specifically, they employ a context-sensitive grammar to accurately model complex dependencies between different parts of an instance. The authors acknowledge that an approach based exclusively on a grammar is not sufficient since the number of proposed graphs grows exponentially with the number of production rules in the grammar. To address the issue, it was proposed to use a Metropolis Hastings (MH) Markov Chain Monte Carlo (MCMC) method, where the sampling problem is reduced to the easier task of simulation. They use context-sensitive graph grammar to inform the MH proposal distribution, but also introduce a probability density estimator to define the MH acceptance procedure. This allows us to deal separately with local and global constraints: the locally context-sensitive graph grammar is used for the local constraints and the regularized statistical model is used for the global or long range constraints. The two approaches complement each other: the grammar is a flexible non-parametric approach that can model complex dependencies between vertices that are within a short path distance from each other; the statistical model, instead, can employ the bias derived from the particular functional family (linear) and the type of regularization used (a penalty over the $\ell_{2}$ norm of the parameters) to generalize long range dependencies to similar cases. This approach is therefore adequate when the underlying concept exhibit local dependencies that are more complex than long range ones. Unfortunately in some application domains instances can exhibit complex long range dependencies between their different constituent parts. This is the case for RNA polymers, long sequences of atomic entities (nucleotides) that self interact, establishing pairwise bounds that can typically span the entire length of the sequence, where sequences can range from hundreds to thousands of atomic parts.
A different way to view the issue of long range dependencies is that of the selection of an appropriate scale of representation at which to work and reason. Certain application domains exhibit natural encodings, i.e. instances are encoded as graphs where nodes represent specific entities, such as nucleotides in the case of RNA sequences. However, it is known that a more effective functional description of RNA polymers can be obtained in terms of structural components such as stems (stretches of consecutive paired nucleotides) and loops (stretches of consecutive unpaired nucleotides). Under this view, dependencies that are local at the coarser scales correspond to longer range dependencies at the original scale.
A constructive system suitable for these domains needs to be able to adequately model complex long range dependencies and is more effective if it operates at a convenient coarser scale rather than that of the individual units. Here we tackle all these issues extending the approach presented in Costa (2017) with two key ideas: 1) we allow a user/domain defined graph coarsening procedure, and 2) we allow domain specific optimization procedures to ensure that generated instances are always viable.
2 Method
Costa (2017) presented an approach to sample graphs from an empirical probability distribution using a variant of the Metropolis Hastings (MH) algorithm (Metropolis et al., 1953). The MH approach decomposes the transition probability of an underlying Markov process in two factors, 1) the proposal and 2) the acceptance-rejection probability distributions. The algorithm starts from a seed element and iteratively proposes new instances that are stochastically accepted if they fit appropriate requirements. The key element for an efficient MH design is building the proposal distribution in such a way that the generated elements will not be rejected too often. To do so Costa suggests to use a graph grammar and to infer its rules from the available data using grammatical inference techniques upgraded to structured domains (i.e. domains where instances are naturally represented as labeled graphs). A graph grammar Rozenberg and Ehrig (1999) is a formal framework used to specify rules for graph transformation. It involves a set of productions denoted as $(M,D,E)$, where $M$ represents the ’mother’ subgraph that is to be replaced, $D$ is the ’daughter’ subgraph that replaces $M$, and $E$ is the embedding mechanism that governs how $D$ is incorporated into the overall graph structure. This formalism allows for the systematic generation of complex graph structures through the iterative application of these transformation rules. Costa (2017) introduced an efficient graph grammar based on the concept of distributional semantics (Harris, 1954, 1968) and on the substitutability principle (Clark and Eyraud, 2007). The notion of mother and daughter graphs are specialized here as neighborhood graphs that share the “outer” context, and the embedding mechanism is simplified by imposing an exact match of the context (see Fig. 1). The context is called interface and its size is parametrized by a discrete value $T$ ($T=2$ means that the “thickness” of the context comprises two nodes), while the inner part is called core and its size is parametrized by the radius $R$. The embedding mechanism can now replace a core-interface pair (CIP) with another, provided that the interfaces are isomorphic. Note that when the radii have different values, the overall size of the generated (daughter) graph is reduced or increased with respect to the size of the seed (mother) graph. One of the advantages of this type of grammar is that the set of all available production rules can be induced efficiently from a set of graphs (see (Costa, 2017) for details). More precisely, a core graph $C^{v}_{R}(G)$, is a neighborhood graph of $G$ of radius $R$ rooted in $v$. An interface graph, denoted $I^{v}_{R,T}(G)$ is the difference graph of two neighborhood graphs of $G$ with the same root in $v$ and radii $R$ and $R+T$. In Fig. 1 the core $C_{R}^{v}(G)$ (dark) is determined by vertices at maximal distance $R$ ($R=0$ in the figure) from a chosen root vertex $v$. The interface $I_{R,T}^{v}(G)$ with thickness $T=2$ is in a lighter shade. Given a new core-interface pair (CIP) with matching interface, the substitution can take place to yield the replacement of a carbon with a nitrogen atom.
2.1 Contribution
In the context of RNA polymers, two key challenges emerge: (1) resolution and (2) viability. First, the level of resolution impacts efficiency, as replacing only a few nodes per iteration is insufficient for structures with hundreds of nodes. Second, viability pertains to the structural integrity of the generated graphs—small alterations can lead to invalid structures. It is known that for RNAs a more effective description is obtained considering larger structural components such as stems and loops and not working at the resolution scale of single nucleotides. As for the second issue, it is known that the function of RNA polymers depends on their global structure (i.e. the set of pairs of interacting nucleotides), which can significantly change when even a single nucleotide is altered. To deal with these issues we propose two enhancements: 1) a grammar coarsening procedure and 2) a constraint integration procedure. The following section elaborates on the details of this approach, demonstrating its effectiveness across two case studies.
Grammar coarsening. The idea is to allow users to specify a coarsening procedure via the notion of edge contraction: an operation which removes an edge from a graph while simultaneously merging the two vertices that it previously joined. In addition, we allow for the more general notion of vertex contraction, which removes the restriction that the contraction must occur over vertices sharing an incident edge. Both procedures are defined using a node attribute $cid\in\mathbb{N}$ called contraction-identifier and contracting all vertices that share the same $cid$. We propose to use the contraction operation as a flexible way to transform a graph to its coarser version, $G\mapsto G^{\prime}$. See Fig. 1 for an example of coarsening. Cores and interfaces can now be defined exploiting the contracted graph. Starting from a CIP on the coarse graph, we define the core as the sub graph induced by the vertices of the original graph that have been contracted to vertices of the core of radius $R$ in the coarse graph, $C_{R}^{v}(G^{\prime},G)$. The new interface graph $I_{R,B}^{v}(G^{\prime},G)$ is defined as the Cartesian product of the graph induced by the nodes adjacent to the core nodes in $G$ at maximal distance $B$ (the thickness in the base graph) and the interface graph on the coarse graph. In words, we require that both the interface at base level and at coarse level match for a core swap to take place. An RNA sequence is a sequence over the nucleotides $A,C,G$ and $U$, these nuceleotides may form hydrogen bonds, forming a structure: $A$ with $U$ and $C$ with $G$. In Fig. 1 we depict an RNA polymer graph at the nucleotide level (center) and its coarse version (right) where the contraction was informed by the notion of structural components such as stems, hairpin loops, and multi-loops. Note that a core at coarse level (in dark) made only of a single node corresponds to multiple nodes at base level. Moreover, while the interface at base level requires only the presence of few nucleotides, these have to belong to a context defined at the coarse level that can span a much larger fraction of the instance and can be viewed as a global localization indicator.
Figure 1: A depiction of core and interface subgraphs. Dark shading indicates the core subgraph, while light shading represents the interface. The diagram on the left shows a molecular graph with its core substitution, while the middle and right diagrams depict RNA encodings at different resolutions (nucleotide-level and structural-element-level, respectively). At nucleotide-level nodes are labeled with nucleotides codes ACGT, while at structural level nodes are labeled as H)airpin, M)ultiloop, S)tem, D)angling end.
Constraint integration. It should be possible to easily make use of specific feasibility constraints when they are available for a given domain. For RNAs the relation between sequence and structure is known to be governed by thermodynamical forces that seek to minimize the amount of free energy of the molecule. The structure of a given sequence can be computed using sophisticated dynamic programming optimization algorithms. To integrate these constraints in the constructive protocol we let the user specify a transformation function which maps candidate instances to feasible instances. In our case, the transformation takes a graph representing an RNA structure constructed by the proposed procedure, removes the pairwise bounds, and recomputes them as the solution of a given folding algorithm. This transformation ensures that we are always computing the acceptance probability on viable RNA structures.
Chemical compounds case. In addition to RNA polymers we consider also small drugs. One type of long range dependency for organic molecules is the membership to ring structures (see Fig. 2). In this case, the context matching procedure is sufficient to guarantee that the CIP replacement operation will almost always generate feasible molecular structures, which eliminates the necessity of a constraint integration procedure.
Figure 2: A molecule and its coarsened version $R=0$ (dark green), $B=1$ (left, lt. green), $T=1$ (right, lt. green). Local constraints are satisfied by the original graph on the left while longer ranging constraints (specifically the fact that the carbon atoms are not only present but form a cycle) is modeled in the coarsened version (right). The coarsening method labeled contracted cycles with the hash of the associated sub graph.
3 Empirical evaluation
Ribonucleic acid (RNA) consists of a chain of nucleotides connected via a ribose backbone. RNA polymers cover essential biological roles ranging from coding, decoding, regulation and expression of genes. Specifially in this work we are looking at non coding RNAs which are not translated into proteins but self interact, forming a structure which is indicative of its function. The RFam database (S. Griffiths-Jones, 2003) groups known RNA sequences in functionally and evolutionarily related families. To empirically investigate the performance of the proposed constructive approach we selected the SAM-I/IV variant riboswitch (RF01725) and SAM riboswitch (S box leader) (RF00162), which are families that exhibit a rich structure (e.g. that do not consist only of a single hairpin element) and an average sequence length of 100. We showcase graphs from RF01725 in Fig.3. The general aim is to synthesize functionally equivalent but novel sequences, a task with important biomedical applications.
Figure 3: Examples of the rf01725 family of RNA, exhibiting a large variety of structural configurations.
When working with RNAs, the feasibility constraints are related to the structure determination. To improve accuracy we use a two step approach: given a candidate sequence we 1) identify $k$ nearest neighbors among the available training instances, then 2) we align the $k+1$ sequences using the MUSCLE (MUltiple Sequence Comparison by Log- Expectation) program (Edgar, 2004) and compute the consensus structure using the RNAalifold program (Stephan H. Bernhart, 2008). This procedure identifies the most representative structure in the ensemble of possible suboptimal solutions. Here we use four neighbors.
In order to evaluate if the constructed sequences are functionally equivalent to the original examples, we use a state-of-the-art classifier as the oracle rather than resorting to a more expensive real world biological experiment. To define the membership of a sequence to a given family, the RFam database uses the covariance model computed by the program Infernal (INFERence of RNA ALignment) (E. P. Nawrocki, 2013) induced over a hand curated set of representative sequences. In Fig. 5 we report the average score achieved by the constructed sequences as the training set increases. The horizontal line indicates the family specific threshold above which the Infernal model accepts a sequence as a valid instance of the family. In blue we report the results for the proposed extension and in red the results for the original method. From the perspective of usual machine learning techniques it seems interesting that the performance decreases when training material is added. This is due to the sampling process. Productions are repeatedly applied to one of the training instances. If the grammar is small, there are only few productions applicable, resulting in a final instance that is close to training instances. To make it comparable to our method, we used MUSCLE and RNAalifold to train a new Infernal model and report the performance of its generated sequences (dashed). We note that the coarsening strategy consistently improves the quality of the results.
In order to evaluate if the generated graphs are distinct from the original material, yet cover a similar distribution, we report an evaluation of the similarity between these two sets. More precisely we compute two measures: 1) the Kullback-Leibler (KL) divergence between the probability density estimated on the original data vs. the density estimated on the generated graphs, and 2) the set similarity between the original and the generated data. For both measures we vectorize the graphs via the NSPDK graph kernel (costagrave) with its default settings. We used the RF01725 RFam family, since the graphs have few nodes and a relatively large structural variety.
Specifically, we compute the symmetrized KL divergence as:
$D_{KL}(P_{G}\parallel P_{R})+D_{KL}(P_{R}\parallel P_{G})$
where $P_{G}$ and $P_{R}$ are the probability distributions estimated from the generated and real data, respectively. To estimate these probabilities, we train two One-Class SVM models: one on the generated data and the other on the real data. We then use these models to evaluate the probability of a test set that consists of an equal mix of real and generated instances. The model trained on the generated data produces the probabilities $P_{G}$, while the model trained on the real data generates $P_{R}$. This procedure is repeated using cross-validation, and we report the median results from these trials.
To measure the similarity between the generated and real graph sets, we employ a normalized kernel similarity measure. Specifically, we compute the kernel similarity between each pair of elements from the set of generated instances $G$ and the set of real instances $R$, summing these values across all pairs:
$K(G,R)=\sum_{g\in G}\sum_{r\in R}k(g,r)$
where $k(g,r)$ is the NSPDK graph kernel, which measures the similarity between an element $g\in G$ (from the set of generated instances) and $r\in R$ (from the set of real instances). To normalize this similarity score, we divide by the geometric mean of the self-similarity scores for both sets:
$\text{similarity}(G,R)=\frac{K(G,R)}{\sqrt{K(G,G)\cdot K(R,R)}}$
For each box in the box plot we collect 7 repetitions.
In Fig. 5 we compute these measures as a function of the training set size. We observe that the similarity tends to decrease as more diverse material becomes available and that also the divergence tends to vanish, indicating that the generated instances induce the same probability density as the original ones, albeit being increasingly different.
Figure 4: Estimated equivalence by Infernal. The horizontal line is the threshold for biological significance.
Figure 5: Similarity to the training instances. The generated graphs should differ from the training set, yet induce the same probability density.
3.1 Graph generative approaches
Deep neural network approaches to generate graphs exist, but they cannot be easily adapted to the domain of large graphs with long range dependencies. DeepGMG (Yujia et al., 2018) and GraphVAE (Simonovsky Martin, 2018) for example can process graphs with a maximum size of few tens of nodes. Domain specific solutions on chemical compounds can operate on larger graphs; the MOSES (Polykovskiy et al., 2020) benchmark contains a large set of graphs, these are restricted to small cycles, an alphabet size of 8 and a molecular weight of 250-350 Daltons (which results in small graphs). Several algorithms, e.g. CharRNN (Segler et al., 2018) which can only operate on SMILES strings, are domain specific.
In many real world cases, the size of the available set of examples is not extremely large. To simulate this scenario we limit the train set size to 1000 instances. We implemented the coarsening as described before and used the parameters $B=2,T=1,R=\{0,1\}$. MOSES offers a variety of measures to assess the quality of the produced molecular graphs: FCD "Frechet ChemNet Distance" a deep learning based estimator that determines how similar instances are to “real” molecules (Preuer et al., 2018); SNN measures how close instances are to the test set on average; Frag and Scaf computes the cosine similarity induced by fragment collections of test and produced sets; IntDiv1,2 measures the internal diversity of the sampled set; filters looks for unstable, very reactive or toxic instances. Furthermore they implement four algorithms for molecule sampling: VAE (Blaschke et al., 2017) a FAN (Sanchez-Lengeling et al., 2017) combines Generative Adversarial Networks (GANs) and reinforcement learning (RL) to work on SMILES, CharRNN emplays a Long Short-Term Memory RNN to learn the language and syntax of SMILES and AAE Adversarial autoencoder, a mix of VAEs and GANs (Polykovskiy et al., 2020).
We examine the drawbacks of each approach. VAE and ORGAN produce less than 10% valid graphs, CharRNN and AAE produce 35% and 36% valid and unique graphs. The best performing approach is CharNN with top performance in FCD, Frag and Scaf measures. The drawback of CharNN is that it produces molecules of such small size (5-7 atoms) that exhaustive enumeration becomes possible. AAE performs best in SNN (this is probably due generating graphs of the appropriate size), but low according to the IntDiv scores(0.69/0.66, while all other generators are in the range 0.84 $\pm$ 0.02), suggesting that its products are very similar to the training material (it outputs only 216 unique graphs over 604 valid ones). The products of the non coarsened grammar perform very similar to the coarsened version however only 79 of the 1000 generated graphs pass the filter test (second lowest is ORGAN with 58.3% , coarsened grammar: 87.2%) this suggests long range dependencies are not met, resulting in unstable molecules. Each weakness of our approach is that the generated instances tend to be large in size which affects the SNN measure. This tendency is a byproduct of the acceptance procedure and in future work will be addressed. In summary, our procedure, while being domain agnostic, generates 100% viable graphs, with comparable performance to the other specialized algorithms (see Tab. 1).
Measure AAE CRNN VAE ORGAN Our
$\uparrow$ valid 0.60 0.351 0.001 0.108 1.0
$\uparrow$ uniq@1K 0.36 1.0 1.0 0.944 1.0
$\downarrow$ FCD/Test 33.9 6.1 nan 33.2 35.8
$\uparrow$ Frag/Test 0.91 0.98 0.0 0.56 0.89
$\uparrow$ Scaf/Test 0.05 0.26 nan 0.0 0.0
$\uparrow$ SNN/Test 0.47 0.37 0.09 0.27 0.22
$\uparrow$ IntDiv 0.69 0.85 0.0 0.85 0.83
$\uparrow$ IntDiv2 0.66 0.84 0.0 0.81 0.82
$\uparrow$ Filters 0.96 0.93 1.0 0.58 0.87
Table 1: FCD - ChemNet distance; a measurement for molecule likeness, Frag/Scaf - cosine similarity to graph fragments, SNN - similarity to the graphs in the test set, IntDiv - diversity of generated artifacts, Filters - fraction of graphs that is non toxic or not unstable. The graph coarsened graph grammar approach performs similar to the specialized methods while generating only valid graphs. CharRNN generates very small graphs, AAE generates graphs that are very similar to those in the training set.
GraphRNN (You et al., 2018) works with unlabeled graphs and sets out to model long range dependencies using an auto regressive model. We Trained this method on the seed sequences of RF01725 whose structure we obtained with the MUSCLE scheme described previously. In Fig. 6 we show three graphs generated by GraphRNN. While the hairpin structure is present, in the lower part of the figure we see cycles and in general more than two dangling ends indicating that the most basic constraint, a continuous ribose backbone, is violated. Out of 1024 generated graphs only 13 passed the following simple filters: a) maximal 2 nodes have a degree of 1. b) none has a degree > 3. c) is connected. d) cycles where exactly 2 nodes have degree 3 and any other vertex has degree 2, the degree 3 nodes must be directly connected. a) and c) are necessary for RNA graphs because an nucleotides form a path graph on top of which H-bonds (edges between non adjacent vertices) are formed. One nucleotide can only form a single H-bond (filter b). Filter d eliminates "nearly-hairpin" structures as seen in Fig. 6(lower left), where the degree 3 nodes next to each other, it would be a valid hairpin structure.
Figure 6: Instances generated by GraphRNN when trained on the seed structures of RF01725. The outputs are not structured enough compared to Fig.3 and there is no clear ribose backbone, violating the domain constraints of RNA.
4 Conclusion
We have introduced an approach to efficiently model long-range dependencies in graph generation, using a coarsening technique. Our method has been tested on the challenging task of RNA molecule generation, which is particularly difficult due to the large graph sizes, limited training data, and the need to preserve complex global structure constraints. The results are promising, demonstrating that our method can generate novel RNA sequences that are functionally equivalent to the original examples. Additionally, we compared our approach to various specialized systems for generating small molecular graphs using deep neural networks, and in scenarios with limited training data, our method achieved comparable performance. Lastly, we found that these approaches are not well-suited for domains with significant long-range dependencies. In future work, one limitation we aim to address is the reliance on predefined coarsening procedures, which may not be optimal for every domain, especially those with limited domain-specific knowledge. To overcome this, we plan to explore ways to learn coarsening procedures in a data-driven fashion, i.e. leveraging machine learning techniques, we hope to dynamically capture these procedures directly from available data. This approach would not only enhance the adaptability of our method across various domains but also improve its scalability in dealing with highly complex and large graph structures. In parallel, we will also investigate novel strategies to reduce the computational costs associated with handling long-range dependencies, which is crucial for the efficient application of our method in modeling real-world biological systems.
5 Funding/Acknowledgements
This work was supported by the de.NBI Cloud within the German Network for Bioinformatics Infrastructure (de.NBI) and ELIXIR-DE (Forschungszentrum Jülich and W-de.NBI-001, W-de.NBI-004, W-de.NBI-008, W-de.NBI-010, W-de.NBI-013, W-de.NBI-014, W-de.NBI-016, W-de.NBI-022).